# Measuring Reactive-Load Impedance with Transmission-Line Resonators Beyond the Perturbative Limit

### v1

@@ I Introduction
Accurate microwave-impedance measurements are essential for probing electronic properties and evaluating material performance in quantum circuits. While capacitance and inductance constitute the imaginary components of the complex impedance, careful analysis of these reactive elements can reveal key insights into electronic structure and correlated states through the quantum capacitance [Xia2009, Riley2015, Maji2024] and kinetic inductance [Kreidel2024, Banerjee2025]. In contrast, the resistive component captures energy dissipation within the material, i.e., the dielectric loss in a capacitor under the equivalent series resistance model. This dielectric loss is a primary factor limiting coherence times in superconducting qubits [Larsen2015, Lange2015, casparis2018, Wang2022, strickland2024, zheng2024], and accurate quantification of such loss is therefore essential for improving qubit performance [Wang2015]. Moreover, minimizing dielectric loss also benefits other superconducting devices, including amplifiers, detectors, and parametric circuits, thereby advancing the full suite of superconducting quantum technologies.
@@
Resonant circuits provide a highly sensitive means of measuring the complex impedance of materials or circuit components. When the device under test is small — such as a lumped element in a quantum circuit or a two-dimensional material with exotic electronic properties — it can be incorporated into a resonator as a load that produces measurable shifts in the resonant frequency $f_{r}$ and internal quality factor $Q_{i}$ [Kreidel2024]. To put this sensitivity in perspective, attaching a $200~\mathrm{fF}$ capacitor with an equivalent series resistance of $200~\mathrm{\mu\Omega}$ to a resonator with $f_{r}=7~\mathrm{GHz}$ and $Q_{i}=10^{6}$ shifts $f_{r}$ by approximately $830~\mathrm{MHz}$ and reduces $Q_{i}$ to $8\times 10^{5}$. The sensitivity of this approach improves further with increasing $Q_{i}$ and greater electromagnetic participation of the load.
@@
Three-dimensional cavities can achieve $Q_{i}>10^{9}$ and support multiple modes [Reagor2013, Lei2023, Checchin2021], but their macroscopic mode volumes limit coupling to nanoscale samples. Planar superconducting resonators overcome this limitation by confining electromagnetic fields near the surface, enabling strong coupling to nanoscale devices while retaining straightforward fabrication and modeling [Krantz2019, Blais2021]. Lumped-element resonators concentrate fields into small volumes and thereby maximize coupling to nanoscale devices, though their nonuniform field distribution complicates loss modeling and parameter extraction. In contrast, transmission-line resonators — typically realized on-chip as coplanar waveguides (CPWs) — combine ease of fabrication, characterization, and modeling with the ability to support multiple harmonic modes [Jin2025, Megrant2012, Kalacheva2023, Antony2021]. However, most analyses of transmission-line resonators rely on perturbative approximations [Gao2008, Kreidel2024, McRae2020], which can limit accuracy, or on numerical simulations [Wang2022, Wenner2011], which demand specialized expertise and significant computational resources for evaluating participation ratios.
@@ II Theoretical Model
Examples of the standing-wave voltage $(V)$ and current $(I)$ profiles in a transmission-line resonator capacitively coupled to a busline are shown in Fig. 1(c). When the coupling capacitance $C_{c}$ is sufficiently small, the resonator exhibits a voltage antinode (and corresponding current node) at one end ($z=0$). The boundary condition is then determined by the termination at the opposite end ($z=$-$\ell$). For $\lvert X\rvert\ll Z_{0}$ (i.e. small inductance or large capacitance), a voltage node forms at the DUT, giving rise to a fundamental $\lambda/4$ resonant mode. Conversely, when $\lvert X\rvert\gg Z_{0}$ (small capacitance or large inductance), a current node forms at the DUT, producing a $\lambda/2$ mode. These limits define perturbative regimes, in which the DUT participation can be treated as a small inductive or capacitive perturbation to an unloaded $\lambda/4$ or $\lambda/2$ resonator, respectively [McRae2020, Kreidel2024, Antony2021].
@@ III Analytic solution and simulation
To calculate the resonant frequency, we solve the transmission-line equations [Pozar] using the appropriate boundary conditions (see Appendix A) and obtain
@@
where $n$ is the standing-wave mode index. This result, consistent with earlier analyses in Ref. [Ramachandran1961], provides a concise expression for the resonance of a transmission-line resonator terminated by an arbitrary reactive load. Although simple in form, it serves as the foundation for thoroughly extracting circuit parameters when extended to include higher-order modes, dissipation, and the concept of the energy participation ratio, as developed in this section.
@@
Here, $Q_{\text{open}}=\pi/(2\alpha l)$, $\alpha$ is the attenuation constant per unit length for the transmission-line resonator. In the limit $\phi\rightarrow n\pi$, these analytic results reduce to the perturbative forms obtained in previous treatments [McRae2020, Kreidel2024, Antony2021]. However, for arbitrary $\phi$, the system can deviate substantially from the perturbative regime, and Eq. 5 provides an explicit analytic relation between the resonator loss and reactive loading that was not available in previous microwave analyses. Together, Eqs. 1, 2, and 5 form a unified formulation valid beyond the perturbative regime, enabling quantitative extraction of reactive and dissipative parameters.
@@
To test the analytic model, we simulate a CPW resonator with $Z_{0}=50~\Omega$ terminated by a series combination of a resistor $r$ and capacitor $C$ using the Qucs circuit simulation software. We set $\alpha$ proportional to $f_{r}$, as the loss is dominated by dielectric loss. The resonance parameters $f_{r}$ and $Q_{i}$ are extracted from the simulated spectra by circle-fit analysis of the $S$-parameters [Probst_circle_fit].
@@
In principle, it is straightforward to determine $\tan\delta$ by measuring $f_{r}$, $f_{\text{open}}$, $Q_{i}$, and $Q_{\text{open}}$, and applying Eqs. 2 and 5. However, statistical variation inherent to resonators limits the accuracy of such extraction. In particular, $Q_{\text{open}}$ exhibits variability not only between nominally identical resonators but also across repeated measurements of the same device [McRae2025]. When uncertainty in $Q_{\text{open}}$ dominants, the fractional uncertainty in $\tan\delta$ is given by:
@@
Similar in spirit to approaches used in 3D superconducting cavities [Lei2023], a multimode self-calibration approach eliminates the need for a separate reference resonator. For each measured mode, Eq. 1 yields the resonant frequency, which, when inserted into Eq. 2, provides the corresponding participation ratio. Together with the measured $Q_{i}$, Eq. 5 then gives one constraint relating $\tan\delta$ and $Q_{\text{open}}$. Measuring at least two resonant modes yields two equations for these two unknowns, enabling simultaneous extraction of $\tan\delta$ and $Q_{\text{open}}$ from a single DUT, without reference to a separate resonator.
The above analysis assumes a known frequency dependence of $Q_{\text{open}}$, which is set by the dominant loss mechanism. In CPW resonators limited by surface or interfacial dielectric loss [Wenner2011, Bruno2015], the attenuation constant $\alpha$ (and thus $Q_{\text{open}}$) increases approximately linearly with frequency [Gao2008, chen2025], as demonstrated in Fig. 6. In contrast, when loss is dominated by weakly conductive dielectrics or resistive elements, $\alpha$ is approximately frequency independent. When $Q_{\text{open}}(f)$ is not known a priori, two modes no longer uniquely determine both $Q_{\text{open}}$ and $\tan\delta$; access to additional modes is required to constrain the loss mechanism. Conversely, once $Q_{\text{open}}(f)$ is established, the same formulations naturally enable extraction of $\tan\delta(f)$ across all accessible modes, if such a frequency dependence exists.
@@ IV Experimental Implementation
To demonstrate the applicability of our explicit closed-form formulations, we present an experimental implementation using fabricated CPW resonators terminated by van der Waals (vdW) parallel-plate capacitors (PPCs) which primarily exhibit capacitive reactance with small, approximately frequency-independent dielectric loss in the microwave regime [Wang2022]. Although demonstrated with capacitive DUTs, the same formulation (Eqs. 12 and 5) applies directly to inductive and hybrid systems. A systematic exploration of these extensions lies beyond the scope of this work but represents a clear direction for future study.
@@
In a parallel-plate capacitor model, the dielectric constant $\kappa_{\text{hBN}}$ is obtained from the slope of a linear fit to the data with the $y$-intercept constrained to zero. Finite-element simulations were used to confirm that the parallel-plate model is applicable, i.e. that fringing fields can be ignored (see Table 3). Frequency errors are negligible compared with geometric uncertainties. The error bars therefore represent uncertainties in the PPC area, estimated to be within 10% due to imperfections (e.g., bubbles) in the van der Waals interface, which can alter the effective capacitor area. From the fits, we find $\kappa_{\text{hBN}}=3.33\pm 0.06$ from the single-mode analysis and $\kappa_{\text{hBN}}=3.06\pm 0.08$ after multimode self-calibration of $f_{\text{open}}$. Both values agree well with the reported out-of-plane dielectric constant of bulk hBN, $\kappa_{\text{hBN}}=3.03$ [Laturia2018]. The closer agreement from multimode calibration arises from the more accurate determination of $f_{\text{open}}$. We find that the multimode-extracted $f_{\text{open}}$ is systematically lower than that obtained from the separate open resonator. This discrepancy may result from a small additional inductance associated with the device geometry and/or a slight modification of the effective resonator length introduced by the Al patch. While the exact origin requires further investigation, neglecting this shift leads to an overestimation of $C_{\text{DUT}}$, and therefore an inflated dielectric constant, in a single-mode analysis.
We further compare the hBN loss tangent obtained from the single-mode analysis with that from the multimode self-calibrated approach; the results are summarized in Tab. 1. In the single-mode method, the loss tangent is obtained by comparing the single-photon $Q_{i}$ of DUT-terminated resonators with the statistical distribution of $Q_{\text{open}}$ measured across nominally identical reference resonators. In contrast, the multimode analysis uses the measured $Q_{i}$ values from the $n=1$ and $n=2$ modes of the same DUT to directly solve for both $\tan\delta$ and $Q_{\text{open}}$, assuming that $Q_{\text{open}}$ is independent of frequency. This assumption is supported by multimode measurements of reference resonators, as shown in Fig. 7.
Fig. 4(c) shows the distribution of $Q_{\text{open}}^{-1}$ for the reference resonators, together with the fundamental-mode $Q_{i}^{-1}$ values of DUTs A–C (vertical dashed lines). The reference resonators exhibit a broad spread in $Q_{\text{open}}$, well described by a log-normal distribution [McRae2025]. Because hBN exhibits low dielectric loss, the DUT $Q_{i}$ values lie near the center of this distribution (around $1.48\times 10^{5}$), where the DUT introduces only a small change to the total resonator loss. As a result, the inferred $\tan\delta$ becomes highly sensitive to device-to-device variability in $Q_{\text{open}}$ rather than to the intrinsic loss of the hBN.
@@ V Conclusion
The analysis further identifies the regime of maximal extraction accuracy, which occurs when the DUT reactance is comparable to the resonator impedance and the DUT energy participation ratio is maximized. We validate the framework through circuit simulations and low-temperature measurements of van der Waals capacitors, demonstrating both quantitative accuracy and experimental feasibility. In addition, the multimode self-calibration approach mitigates uncertainties arising from reference-resonator variability by solving for both $f_{\text{open}}$ and $Q_{\text{open}}$ within a single device. Beyond improving accuracy, this multimode analysis provides physical discrimination between capacitive and inductive components of the DUT reactance.
@@
ACKNOWLEDGEMENTS This work was primarily supported by the Army Research Office under Contract W911NF-22-C-0021 (NextNEQST SuperVan-2). Synthesis of boron nitride (K.W. and T.T.) was supported by the Elemental Strategy Initiative conducted by the MEXT, Japan (Grant No. JPMXP0112101001) and JSPS KAKENHI (Grant Nos. JP19H05790 and JP20H00354). J.H. acknowledges support from the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF10277. J.P. acknowledges support from the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00358841) and the education and training program of the Quantum Information Research Support Center, funded through the NRF by the MSIT of the Korean government (No. 2021M3H3A1036573). J.B. acknowledges support from the Army Research Office under Grant Number W911NF-24-1-0133. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix D Device fabrication
To define the electrical contacts, we spin-coat the chip with a bilayer of PMMA A4/MMA EL10 resist and pattern the contact regions using electron-beam lithography. These contacts connect one side of the NbSe2 flake to the resonator center conductor and the other side to the ground plane. Before metal deposition, we perform in situ Ar ion milling to remove surface oxides from both the NbSe2 and Nb contact regions. Finally, aluminum (Al) is deposited to form high-transparency superconducting contact between the vdW material and the Nb resonator [Antony2021].

### v2

@@ I Introduction
Accurate microwave-impedance measurements are essential for probing electronic properties and evaluating material performance in quantum circuits. While capacitance and inductance constitute the imaginary components of the complex impedance, careful analysis of these reactive elements can reveal key insights into electronic structure and correlated states through the quantum capacitance [30, 24, 17] and kinetic inductance [13, 2]. In contrast, the resistive component captures energy dissipation within the material, i.e., the dielectric loss in a capacitor under the equivalent series resistance model. This dielectric loss is a primary factor limiting coherence times in superconducting qubits [14, 8, 5, 28, 25, 31], and accurate quantification of such loss is therefore essential for improving qubit performance [27]. Moreover, minimizing dielectric loss also benefits other superconducting devices, including amplifiers, detectors, and parametric circuits, thereby advancing the full suite of superconducting quantum technologies.
@@
Resonant circuits provide a highly sensitive means of measuring the complex impedance of materials or circuit components. When the device under test is small — such as a lumped element in a quantum circuit or a two-dimensional material with exotic electronic properties — it can be incorporated into a resonator as a load that produces measurable shifts in the resonant frequency $f_{r}$ and internal quality factor $Q_{i}$ [13]. To put this sensitivity in perspective, attaching a $200~\mathrm{fF}$ capacitor with an equivalent series resistance of $200~\mathrm{\mu\Omega}$ to a resonator with $f_{r}=7~\mathrm{GHz}$ and $Q_{i}=10^{6}$ shifts $f_{r}$ by approximately $830~\mathrm{MHz}$ and reduces $Q_{i}$ to $8\times 10^{5}$. The sensitivity of this approach improves further with increasing $Q_{i}$ and greater electromagnetic participation of the load.
@@
Three-dimensional cavities can achieve $Q_{i}>10^{9}$ and support multiple modes [23, 16, 6], but their macroscopic mode volumes limit coupling to nanoscale samples. Planar superconducting resonators overcome this limitation by confining electromagnetic fields near the surface, enabling strong coupling to nanoscale devices while retaining straightforward fabrication and modeling [12, 3]. Lumped-element resonators concentrate fields into small volumes and thereby maximize coupling to nanoscale devices, though their nonuniform field distribution complicates loss modeling and parameter extraction. In contrast, transmission-line resonators — typically realized on-chip as coplanar waveguides (CPWs) — combine ease of fabrication, characterization, and modeling with the ability to support multiple harmonic modes [10, 19, 11, 1]. However, most analyses of transmission-line resonators rely on perturbative approximations [9, 13, 18], which can limit accuracy, or on numerical simulations [28, 29], which demand specialized expertise and significant computational resources for evaluating participation ratios.
@@ II Theoretical Model
Examples of the standing-wave voltage $(V)$ and current $(I)$ profiles in a transmission-line resonator capacitively coupled to a busline are shown in Fig. 1(c). When the coupling capacitance $C_{c}$ is sufficiently small, the resonator exhibits a voltage antinode (and corresponding current node) at one end ($z=0$). The boundary condition is then determined by the termination at the opposite end ($z=$-$\ell$). For $\lvert X\rvert\ll Z_{0}$ (i.e. small inductance or large capacitance), a voltage node forms at the DUT, giving rise to a fundamental $\lambda/4$ resonant mode. Conversely, when $\lvert X\rvert\gg Z_{0}$ (small capacitance or large inductance), a current node forms at the DUT, producing a $\lambda/2$ mode. These limits define perturbative regimes, in which the DUT participation can be treated as a small inductive or capacitive perturbation to an unloaded $\lambda/4$ or $\lambda/2$ resonator, respectively [18, 13, 1].
@@ III Analytic solution and simulation
To calculate the resonant frequency, we solve the transmission-line equations [20] using the appropriate boundary conditions (see Appendix A) and obtain
@@
where $n$ is the standing-wave mode index. This result, consistent with earlier analyses in Ref. [22], provides a concise expression for the resonance of a transmission-line resonator terminated by an arbitrary reactive load. Although simple in form, it serves as the foundation for thoroughly extracting circuit parameters when extended to include higher-order modes, dissipation, and the concept of the energy participation ratio, as developed in this section.
@@
Here, $Q_{\text{open}}=\pi/(2\alpha l)$, $\alpha$ is the attenuation constant per unit length for the transmission-line resonator. In the limit $\phi\rightarrow n\pi$, these analytic results reduce to the perturbative forms obtained in previous treatments [18, 13, 1]. However, for arbitrary $\phi$, the system can deviate substantially from the perturbative regime, and Eq. 5 provides an explicit analytic relation between the resonator loss and reactive loading that was not available in previous microwave analyses. Together, Eqs. 1, 2, and 5 form a unified formulation valid beyond the perturbative regime, enabling quantitative extraction of reactive and dissipative parameters.
@@
To test the analytic model, we simulate a CPW resonator with $Z_{0}=50~\Omega$ terminated by a series combination of a resistor $r$ and capacitor $C$ using the Qucs circuit simulation software. We set $\alpha$ proportional to $f_{r}$, as the loss is dominated by dielectric loss. The resonance parameters $f_{r}$ and $Q_{i}$ are extracted from the simulated spectra by circle-fit analysis of the $S$-parameters [21].
@@
In principle, it is straightforward to determine $\tan\delta$ by measuring $f_{r}$, $f_{\text{open}}$, $Q_{i}$, and $Q_{\text{open}}$, and applying Eqs. 2 and 5. However, statistical variation inherent to resonators limits the accuracy of such extraction. In particular, $Q_{\text{open}}$ exhibits variability not only between nominally identical resonators but also across repeated measurements of the same device [26]. When uncertainty in $Q_{\text{open}}$ dominants, the fractional uncertainty in $\tan\delta$ is given by:
@@
Similar in spirit to approaches used in 3D superconducting cavities [16], a multimode self-calibration approach eliminates the need for a separate reference resonator. For each measured mode, Eq. 1 yields the resonant frequency, which, when inserted into Eq. 2, provides the corresponding participation ratio. Together with the measured $Q_{i}$, Eq. 5 then gives one constraint relating $\tan\delta$ and $Q_{\text{open}}$. Measuring at least two resonant modes yields two equations for these two unknowns, enabling simultaneous extraction of $\tan\delta$ and $Q_{\text{open}}$ from a single DUT, without reference to a separate resonator.
The above analysis assumes a known frequency dependence of $Q_{\text{open}}$, which is set by the dominant loss mechanism. In CPW resonators limited by surface or interfacial dielectric loss [29, 4], the attenuation constant $\alpha$ increases approximately linearly with frequency [9, 7], which should yield a frequency-independent internal quality factor. As demonstrated in Fig. 7, our reference resonators exhibit comparable internal quality factors at the fundamental and first harmonic modes, justifying the use of a frequency-dependent $\alpha$ in our simulation. In contrast, when loss is dominated by weakly conductive dielectrics or resistive elements, $\alpha$ is approximately frequency independent. When $Q_{\text{open}}(f)$ is not known a priori, two modes no longer uniquely determine both $Q_{\text{open}}$ and $\tan\delta$; access to additional modes is required to constrain the loss mechanism. Conversely, once $Q_{\text{open}}(f)$ is established, the same formulations naturally enable extraction of $\tan\delta(f)$ across all accessible modes, if such a frequency dependence exists.
@@ IV Experimental Implementation
To demonstrate the applicability of our explicit closed-form formulations, we present an experimental implementation using fabricated CPW resonators terminated by van der Waals (vdW) parallel-plate capacitors (PPCs) which primarily exhibit capacitive reactance with small, approximately frequency-independent dielectric loss in the microwave regime [28]. Although demonstrated with capacitive DUTs, the same formulation (Eqs. 12 and 5) applies directly to inductive and hybrid systems. A systematic exploration of these extensions lies beyond the scope of this work but represents a clear direction for future study.
@@
In a parallel-plate capacitor model, the dielectric constant $\kappa_{\text{hBN}}$ is obtained from the slope of a linear fit to the data with the $y$-intercept constrained to zero. Finite-element simulations were used to confirm that the parallel-plate model is applicable, i.e. that fringing fields can be ignored (see Table 3). Frequency errors are negligible compared with geometric uncertainties. The error bars therefore represent uncertainties in the PPC area, estimated to be within 10% due to imperfections (e.g., bubbles) in the van der Waals interface, which can alter the effective capacitor area. From the fits, we find $\kappa_{\text{hBN}}=3.33\pm 0.06$ from the single-mode analysis and $\kappa_{\text{hBN}}=3.06\pm 0.08$ after multimode self-calibration of $f_{\text{open}}$. Both values agree well with the reported out-of-plane dielectric constant of bulk hBN, $\kappa_{\text{hBN}}=3.03$ [15]. The closer agreement from multimode calibration arises from the more accurate determination of $f_{\text{open}}$. We find that the multimode-extracted $f_{\text{open}}$ is systematically lower than that obtained from the separate open resonator. This discrepancy may result from a small additional inductance associated with the device geometry and/or a slight modification of the effective resonator length introduced by the Al patch. While the exact origin requires further investigation, neglecting this shift leads to an overestimation of $C_{\text{DUT}}$, and therefore an inflated dielectric constant, in a single-mode analysis.
We further compare the hBN loss tangent obtained from the single-mode analysis with that from the multimode self-calibrated approach; the results are summarized in Tab. 1. In the single-mode method, the loss tangent is obtained by comparing the single-photon $Q_{i}$ of DUT-terminated resonators with the statistical distribution of $Q_{\text{open}}$ measured across nominally identical reference resonators. In contrast, the multimode analysis uses the measured $Q_{i}$ values from the $n=1$ and $n=2$ modes of the same DUT to directly solve for both $\tan\delta$ and $Q_{\text{open}}$, again noting that $Q_{\text{open}}$ is independent of frequency.
Fig. 4(c) shows the distribution of $Q_{\text{open}}^{-1}$ for the reference resonators, together with the fundamental-mode $Q_{i}^{-1}$ values of DUTs A–C (vertical dashed lines). The reference resonators exhibit a broad spread in $Q_{\text{open}}$, well described by a log-normal distribution [26]. Because hBN exhibits low dielectric loss, the DUT $Q_{i}$ values lie near the center of this distribution (around $1.48\times 10^{5}$), where the DUT introduces only a small change to the total resonator loss. As a result, the inferred $\tan\delta$ becomes highly sensitive to device-to-device variability in $Q_{\text{open}}$ rather than to the intrinsic loss of the hBN.
@@ V Conclusion
The analysis further identifies the regime of maximal extraction accuracy, which occurs when the DUT reactance is comparable to the resonator impedance and the DUT energy participation ratio is maximized. We validate the framework through circuit simulations and low-temperature measurements of van der Waals capacitors, demonstrating both quantitative accuracy and experimental feasibility. In addition, the multimode self-calibration approach mitigates uncertainties arising from reference-resonator variability by solving for both $f_{\text{open}}$ and $Q_{\text{open}}$ within a single device. Beyond improving accuracy, this multimode analysis provides physical discrimination between capacitive and inductive components of the DUT reactance. Once the effective DUT reactance is experimentally determined, electromagnetic simulations can be used, if desired, to relate it to specific device geometry and material properties.
@@
ACKNOWLEDGEMENTS This work was primarily supported by the Army Research Office under Contract W911NF-22-C-0021 (NextNEQST SuperVan-2). Synthesis of boron nitride (K.W. and T.T.) was supported by the Elemental Strategy Initiative conducted by the MEXT, Japan (Grant No. JPMXP0112101001) and JSPS KAKENHI (Grant Nos. JP19H05790 and JP20H00354). J.H. acknowledges support from the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF10277. J.P. acknowledges support from the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00358841, RS-2024-00393599, IITP-2025-RS-2022-00164799) and the education and training program of the Quantum Information Research Support Center, funded through the NRF by the MSIT of the Korean government (No. 2021M3H3A1036573). J.B. acknowledges support from the Army Research Office under Grant Number W911NF-24-1-0133. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix D Device fabrication
To define the electrical contacts, we spin-coat the chip with a bilayer of PMMA A4/MMA EL10 resist and pattern the contact regions using electron-beam lithography. These contacts connect one side of the NbSe2 flake to the resonator center conductor and the other side to the ground plane. Before metal deposition, we perform in situ Ar ion milling to remove surface oxides from both the NbSe2 and Nb contact regions. Finally, aluminum (Al) is deposited to form high-transparency superconducting contact between the vdW material and the Nb resonator [1].